OpenStudy (anonymous):

In a college biology class, only half of the class earned a passing grade. Out of all the students, 45% studied and passed. What is the probability of someone having studied, given that they passed? A. 1% B. 55% C. 90% D. 45%

OpenStudy (anonymous):

I think "D"

OpenStudy (anonymous):

C. 90%

OpenStudy (anonymous):

who u get that?

OpenStudy (anonymous):

P(having studies given that they passed) = P(studied and pass) / P(passed) = 0.45/ 0.5 = 0.9

OpenStudy (anonymous):

oh ok